1. The wavelength of a particular color of yellow light is 595 nm.


The energy of this wavelength of light is  kJ/photon.

keep in mind sig figs


2. Gamma ray radiation has wavelengths from 1.0×10-16 to 1.0×10-11 m, whereas the wavelength region for radio wave radiation is 10 to 1000 m.  

We can say that:

1. The frequency of gamma ray radiation is _________either higher than lower than the same as radio wave radiation.

2. The speed of gamma ray radiation is _________either higher than lower than the same as radio wave radiation.
Answer #1

1 . E = hv , where E = energy , h= Planck's constant and v = frequency which is equal to c/lambda , where C= speed of light and lambda is wavelength .

So , E = h × c/lambda .

E = (6.626 × 10^-34 joules/s )× (3.0×10^8 m/s) /595×10^-9 m

E = 3.34 × 10^-19 joules = 3.34 × 10^-16 kJ/mole .

In terms of kJ/photon ,

E = 3.34×10^-16 /6.023×10^23 = 5.55 ×10^-40 kJ/photon .